OpenStudy (anonymous):

Silicon tetrafluoride gas can be produced by the following reaction: SiO2+4HF=SiF4+2H2O If 1.00 L of HF gas under pressure at 3.48atm and a temperature of 25 degrees celcius reacts completely with SiO2 to form SiF4. What volume of SiF4 at 15 degrees celcius and .940atm is produced?

OpenStudy (anonymous):

First, use the ideal gas equation to find the number of HF in mol. Be sure to convert the temperature to the Kelvin scale. here HF is the limiting reagent and according to the stoichiometry the number of SiF4 formed is a quarter of the number of HF. Now that we have the number of SiF4. Assuming it's in gas phase, we use the ideal gas equation to find the volume because both P and T are given.

OpenStudy (anonymous):

Give me some time to crunch the numbers on the calculator.

OpenStudy (anonymous):

So the number of HF = P V / (R T ) = (3.48 * 1 ) / ( 0.0821 * 298.15 ) = 0.142 moles. Note that the units are atm for pressure, liter for volume, kelvin for temperature, and L atm per mol per kelvin for temperature the number of SiF4 = 0.25 * number of HF = .0355 moles Now we get the volume of SiF4 with the ideal gas law: volume = (the number of SiF4) R T / pressure = ( .0355 * 0.0821 * 288.15 ) / .940 = .89 liter

OpenStudy (anonymous):

the answer has two sig figs assuming that the "15 degrees" you mentioned is the number "15." which has 2 sig figs.